OpenStudy (anonymous):

Solve the sysyem by substitution. 4x-7y=15 x-3y=5 The answers are (2,-1) Help?

ganeshie8 (ganeshie8):

from the second equation, x-3y = 5 x = 5 + 3y

ganeshie8 (ganeshie8):

substitute this x value in first equation

ganeshie8 (ganeshie8):

4(5+3y) - 7y = 15

ganeshie8 (ganeshie8):

you can solve above for y ?

OpenStudy (anonymous):

I keep coming up with y=-1

ganeshie8 (ganeshie8):

thats the answer says also (: good work :)

OpenStudy (anonymous):

Oh...never mind. That is the y

OpenStudy (anonymous):

And then sub the y and solve for x

ganeshie8 (ganeshie8):

yup soo many subs... thats why its called subb method !!

OpenStudy (anonymous):

So my main problem is there is so many way to do this and sometimes the fractions get so ugly

OpenStudy (anonymous):

I would rather just use elimination...so much easier

ganeshie8 (ganeshie8):

yes i agree.... i prefer elimination most of the times... substitution i feel long process lol

OpenStudy (anonymous):

Like taking 4x-7y=15 -7y=-4x+15 y=4/3x-15/7 And so on, and so on.... It just gets so jumbled

OpenStudy (anonymous):

Anyway...thank you very much!!

OpenStudy (anonymous):

I meant y=-4/7-15/7 :)

ganeshie8 (ganeshie8):

you took the deadly poison... substituting y=-1 in 2nd equation is somewhat simple : x-3y=5 x-3(-1) = 5 x + 3 = 5 x = 2 ;)